Find the force of gravity between you(70.0kg) and a pencil (1.20g) at a distance of 0.500m
1
Expert's answer
2011-06-07T06:30:17-0400
The force of gravity can be calculated by the following formula:

F = (G m1 m2) / r2 = (6.673×10-11* 70* 1.2) / 0.52 = 2.2×10-8 N.